OpenStudy (anonymous):

How do you solve dy/dt = t/(y+1)? Don't you get 2 y's on one side of the equation, so how would you solve for y(t)?

OpenStudy (anonymous):

1. multiply both sides by dt, you get dy = (t dt) /(y+1) 2. multiply both sides by (y+1), you get (y+1) dy = t dt 3. integrate both sides, you get (y^2)/2 + y = (t^2)/2, which is the implicit form of the solution